OpenStudy (anonymous):

3|2x + 2| − 7 = 5 for x. Show your work please.

OpenStudy (anonymous):

3|2x+2| - 7 = 5 3|2x+2| = 5 + 7 3|2x+2| = 12 |2x+2| = 12/3 |2x+2| = 4 2x+2 = 4 or 2x+2 = -4 2x = 4 -2 or 2x = -4 - 2 2x = 2 or 2x = -6 x = 2/2 or x = -6/3 x =1 or x = -2 ----------These are the solutions for x.